OpenStudy (anonymous):

Given the linear equation y= 1/3x-3, find the y-coordinates of the points (-6, ), (-3, ), and (6, ). Please show all of your work

OpenStudy (anonymous):

y= (1/3x) - 3 Put x = -6 y = (1/3)*(-6) - 3 = -2-3 = -5 Point will be (-6, -5). Put x = -3 y = (1/3)*(-3) - 3 = -1-3 = -4 Point will be (-3, -4). Put x = 6 y = (1/3)*(6) - 3 = 2-3 = -1 Point will be (6, -1).